# Search for new physics with charm rare decays at BESIII

### v1

@@ 3 Searches for FCNC processes
FCNC transitions such as $c\to u$ manifest solely through loop-level interactions due to Glashow-Iliopoulos-Maiani (GIM) mechanism [27], which result in low branching fractions (BFs) of the corresponding decays. The FCNC decay $J/\psi\to D^{0}l^{+}l^{-}$ is expected to have a BF on the order of $10^{-13}$ in the SM [28]. In comparison, the FCNC process $J/\psi\to D^{0}\gamma$ is expected with a 1 or 2 orders of magnitude larger BF, due to the presence of one fewer decay vertex. Nevertheless, various new physics (NP) models beyond the SM, such as the Top-Color model [29], the supersymmetric extensions of the SM with or without R-parity violation [30], and the two-Higgs doublet model [31], suggest these BFs may be enhanced to exceed the SM prediction. Any enhancement of the BF with respect to the SM would be a strong indication of NP.
@@ 3.1 Searches for $J/\psi\to D^{0}\mu\mu$ and $J/\psi\to D^{0}\gamma$
For $J/\psi\to D^{0}\mu^{+}\mu^{-}$, no significant signal are observed, and the upper limits on BFs are set to be $\mathcal{B}(J/\psi\to D^{0}\mu^{+}\mu^{-})=1.1\times 10^{-7}$ at the 90% C.L. after considering all systematic uncertainties. This is the first search for a charmonium FCNC process with muons in the final state. This result is compatible with the SM prediction on its BF of $10^{-13}$.
For $J/\psi\to D^{0}\gamma$, no significant signal are observed, and the upper limits on BFs are set to be $\mathcal{B}(J/\psi\to D^{0}\gamma)=9.1\times 10^{-8}$ at the 90% C.L. after considering all systematic uncertainties. Although our measurement does not yet reach the precision predicted by the SM, it provides a valuable reference for probing various new-physics models and restricting their parameter space.
Figure 2: The distributions of $M_{D^{0}\mu^{+}\mu^{-}}$ for $J/\psi\to D^{0}\mu^{+}\mu^{-}$ of the selected candidates in data, signal MC sample, and inclusive MC sample. The simultaneous fit result is shown in the top left, and the individual fit result for each mode is shown in the other three sub-figures, respectively. The black dots with error bars are data, the magenta dotted-dashed line shows the shape of signal MC sample scaled to $\mathcal{B}(J/\psi\to D^{0}\mu^{+}\mu^{-})=1.0\times 10^{-6}$ for the three tags. The blue shaded histogram is the inclusive MC sample, the red line is the fit result, and the blue solid lines is the fitted background.
Figure 3: The unbinned simultaneous fit on the invariant mass distributions for $J/\psi\to D^{0}\gamma$ of modes (a) I, (b) II and (c) III, where the dots with error bar are data, the blue lines are the fit result, and the red dotted and purple curves are the fitted signal and background, respectively.
@@ 4 Searches for $J/\psi$ weak decays
Via the weak interaction, the $J/\psi$ can potentially decay into a single charm meson via such as D accompanied by some non-charm mesons, some of them have been searched for by BESIII[33, 34, 35, 36, 37, 38, 40, 39]. The inclusive BF of charmonium rare weak decays is predicted to be of the order of $10^{-8}$ in the SM [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. If a signal for one of these decays is observed with BFs in the range of $10^{-8}$ to $10^{-6}$, it would indicate new physics beyond the SM.
@@ 4.1 Search for $J/\psi$ weak decays containing a D meson
Figure 6: Fits of the accepted candidates to the recoiling mass spectra for (a) $J/\psi\to\bar{D}^{0}\pi^{0}$, (b) $J/\psi\to\bar{D}^{0}\eta$, (c) $J/\psi\to\bar{D}^{0}\rho^{0}$, (d) $J/\psi\to D^{-}\pi^{+}$, and (e) $J/\psi\to D^{-}\rho^{+}$. The dots with error bars are data and the orange dotted lines are polynomial functions describing the background. The blue solid curves are the total fits. The inclusive MC samples are shown by the green filled histograms.
@@ 4.2 Search for $J/\psi$ weak decays
Full reconstruction of $D_{s}$ mesons with non-leptonic-decay modes does not offer good sensitivity due to the large hadronic background from $J/\psi$ inclusive decays. Therefore, $D_{s}$ candidates are reconstructed via the semi-leptonic decay $D_{s}^{-}\to\phi e^{-}\bar{\nu}_{e}$ with $\phi\to K^{+}K^{-}$. For the $\rho^{+}$ side, the subsequent signal decays are $\rho^{+}\to\pi^{+}\pi^{0}$, $\pi^{0}\to\gamma\gamma$. Unbinned extended maximum likelihood fits are performed to the distributions of the mass recoiling against the $\pi^{+}$ or $\rho^{+}$, denotes as $M_{D_{s}}$, to determine the signal yields, as shown in Fig. 7. No significant signal is observed and ULs on their BFs are set at $\mathcal{B}(J/\psi\to D_{s}^{-}\rho^{+})<8.0\times 10^{-7}$ and $\mathcal{B}(J/\psi\to D_{s}^{-}\pi^{+})<4.1\times 10^{-7}$ at the 90% C.L. In comparison to the previous best limits, the UL for $J/\psi\to D_{s}^{-}\rho^{+}$ has been improved by about an order of magnitude, and the UL for $J/\psi\to D_{s}^{-}\pi^{+}$ has been improved by about three orders of magnitude.

### v2

@@ 3 Searches for FCNC processes
FCNC transitions such as $c\to u$ manifest solely through loop-level interactions due to the Glashow-Iliopoulos-Maiani (GIM) mechanism [27], which result in low branching fractions (BFs) of the corresponding decays. The FCNC decay $J/\psi\to D^{0}l^{+}l^{-}$ is expected to have a BF on the order of $10^{-13}$ in the SM [28]. In comparison, the FCNC process $J/\psi\to D^{0}\gamma$ is expected with a 1 or 2 orders of magnitude larger BF, due to the presence of one fewer decay vertex. Nevertheless, various new physics (NP) models beyond the SM, such as the Top-Color model [29], the supersymmetric extensions of the SM with or without R-parity violation [30], and the two-Higgs doublet model [31], suggest these BFs may be enhanced to exceed the SM prediction. Any enhancement of the BF with respect to the SM would be a strong indication of NP.
@@ 3.1 Searches for $J/\psi\to D^{0}\mu\mu$ and $J/\psi\to D^{0}\gamma$
For $J/\psi\to D^{0}\mu^{+}\mu^{-}$, no significant signal is observed, and the upper limits on BFs are set to be $\mathcal{B}(J/\psi\to D^{0}\mu^{+}\mu^{-})=1.1\times 10^{-7}$ at the 90% C.L. after considering all systematic uncertainties. This is the first search for a charmonium FCNC process with muons in the final state. This result is compatible with the SM prediction on its BF of $10^{-13}$.
For $J/\psi\to D^{0}\gamma$, no significant signal is observed, and the upper limits on BFs are set to be $\mathcal{B}(J/\psi\to D^{0}\gamma)=9.1\times 10^{-8}$ at the 90% C.L. after considering all systematic uncertainties. Although our measurement does not yet reach the precision predicted by the SM, it provides a valuable reference for probing various new-physics models and restricting their parameter space.
Figure 2: The distributions of $M_{D^{0}\mu^{+}\mu^{-}}$ for $J/\psi\to D^{0}\mu^{+}\mu^{-}$ of the selected candidates in data, signal MC sample, and inclusive MC sample. The simultaneous fit result is shown in the top left, and the individual fit result for each mode is shown in the other three sub-figures, respectively. The black dots with error bars are data, the magenta dotted-dashed line shows the shape of the signal MC sample scaled to $\mathcal{B}(J/\psi\to D^{0}\mu^{+}\mu^{-})=1.0\times 10^{-6}$ for the three tags. The blue shaded histogram is the inclusive MC sample, the red line is the fit result, and the blue solid lines are the fitted background.
Figure 3: The unbinned simultaneous fit on the invariant mass distributions for $J/\psi\to D^{0}\gamma$ of modes (a) I, (b) II, and (c) III, where the dots with error bar are data, the blue lines are the fit result, and the red dotted and purple curves are the fitted signal and background, respectively.
@@ 4 Searches for $J/\psi$ weak decays
Via the weak interaction, the $J/\psi$ can potentially decay into a single charm meson via such as D, accompanied by some non-charm mesons. Some of them have been searched for by BESIII[33, 34, 35, 36, 37, 38, 40, 39]. The inclusive BF of charmonium rare weak decays is predicted to be of the order of $10^{-8}$ in the SM [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. If a signal for one of these decays is observed with BFs in the range of $10^{-8}$ to $10^{-6}$, it would indicate new physics beyond the SM.
@@ 4.1 Search for $J/\psi$ weak decays containing a D meson
Figure 6: Fits of the accepted candidates to the recoiling mass spectra for (a) $J/\psi\to\bar{D}^{0}\pi^{0}$, (b) $J/\psi\to\bar{D}^{0}\eta$, (c) $J/\psi\to\bar{D}^{0}\rho^{0}$, (d) $J/\psi\to D^{-}\pi^{+}$, and (e) $J/\psi\to D^{-}\rho^{+}$. The dots with error bars are data, and the orange dotted lines are polynomial functions describing the background. The blue solid curves are the total fits. The inclusive MC samples are shown by the green filled histograms.
@@ 4.2 Search for $J/\psi$ weak decays
Full reconstruction of $D_{s}$ mesons with non-leptonic-decay modes does not offer good sensitivity due to the large hadronic background from $J/\psi$ inclusive decays. Therefore, $D_{s}$ candidates are reconstructed via the semi-leptonic decay $D_{s}^{-}\to\phi e^{-}\bar{\nu}_{e}$ with $\phi\to K^{+}K^{-}$. For the $\rho^{+}$ side, the subsequent signal decays are $\rho^{+}\to\pi^{+}\pi^{0}$, $\pi^{0}\to\gamma\gamma$. Unbinned extended maximum likelihood fits are performed to the distributions of the mass recoiling against the $\pi^{+}$ or $\rho^{+}$, denoted as $M_{D_{s}}$, to determine the signal yields, as shown in Fig. 7. No significant signal is observed and ULs on their BFs are set at $\mathcal{B}(J/\psi\to D_{s}^{-}\rho^{+})<8.0\times 10^{-7}$ and $\mathcal{B}(J/\psi\to D_{s}^{-}\pi^{+})<4.1\times 10^{-7}$ at the 90% C.L. In comparison to the previous best limits, the UL for $J/\psi\to D_{s}^{-}\rho^{+}$ has been improved by about an order of magnitude, and the UL for $J/\psi\to D_{s}^{-}\pi^{+}$ has been improved by about three orders of magnitude.